OpenStudy (anonymous):

the sum of 51 5/6 + (-37 1/12 ) is equal to what?

OpenStudy (jack1):

5/6 = 10/12 so 51 5/6 - 37 1/12 = 51 10/12 - 37 1/12 now subtract, starting with the fractions 10/12 - 1/12 = ? then 51 - 37 = ?

OpenStudy (anonymous):

9/12 and 14

OpenStudy (anonymous):

ok i got it! i think.... is it 14 3/4?

OpenStudy (jack1):

yep, perfect @emdrewer

OpenStudy (anonymous):

thanks! :)

OpenStudy (jack1):

np, hava gud one ;)